OpenStudy (anonymous):

What is the negative square root of 585 simplified to?

OpenStudy (campbell_st):

you need the identity of \[i^2 = -1\] so you are now looking at \[\sqrt{-585} = \sqrt{585i^2}\] now to simplify 585 you have 585 = 9 times 65 so \[\sqrt{-585} = \sqrt{9 \times 65 \times i^2} = 3i \sqrt{65}\] hope it helps

OpenStudy (anonymous):

how do you put that in standard form?